Hint Chec 29> A highway has an optional toll lane that drivers may take to reduce Number of miles traveled on toll lane Toll ($) the time they spend driving. Drivers pay a small fee to enter the toll lane ($0.25). Then, once they leave the toll lane, they pay a fee based on the number of miles they have traveled on the toll lane. Assume that the driver may leave the lane after any whole number of miles and pays for exactly that number, without rounding up. Note that there is a linear relationship between the number of miles a vehicle has traveled and the price of the toll. A. If Frank is on the toll road for 6.00 miles and then leaves the lane, how much will he have to pay total for the trip? 0 0.25 1.00 1.75 4.00 7.75 10 B. Each day, Susan has to pay a toll of $9.25 when she uses the toll lane to get to school. How many miles does Susan travel on the toll lane to get to school? mi about us careers privacy policy terms of use contact us help
Answer #1

Toll entry fee = $ 0.25

A. Frank travels 6 miles then leaves the lane.

Frank has to pay = Entry Fee + toll charges for number of miles travelled

Frank pay = $ 0.25(entry fee) + $4 (for 5 miles) + $1 (for 1 mile)

Frank pay = $ 5.25

B. Susan pay everyday = $ 9.25

Susan pay $ 9.25 for travelling in the toll lane.

Susan first travel 10 miles so pays = 0.25 + 7.75 = $ 8

Next she travels 1 mile so pays = 0.25 + 1 = $ 1.25

Susan total payment = $ 8 + $1.25 = $ 9.25

Susan travels 11 miles.

Assume that every time you enter the lane you have to pay entry fee.